OpenStudy (anonymous):

which is the correct relationship between current , voltage and power ? a decrease in power will occur if current is decreased and voltage remains the same an increase in voltage with no change in current will result in an increase in power in order to increase power there must be a decrease in either current voltage the only way to decrease current is to decrease power

OpenStudy (anonymous):

please help i think its b

OpenStudy (anonymous):

yes this can be a relationship when we are taking about KVA or VA powers.

OpenStudy (rock_mit182):

P = V * I OR: 1. IF V = K(constant) --> I = P/K ; 2. IF I = C (constant) --> V = P/C

OpenStudy (anonymous):

so its c

OpenStudy (rock_mit182):

well if you have a transformer ( alternating current). you have, P = no varies

OpenStudy (rock_mit182):

so I would say D. I hope someone confirm that fact . .

OpenStudy (radar):

Stay with your original thoughts. @hotdog97